OpenStudy (anonymous):

solve the equations, x^2 + 6x = 0

OpenStudy (anonymous):

x=-6,0

OpenStudy (callisto):

x^2 + 6x = 0 x(x+6) =0 x=0 or x+6 =0 x=0 or x=-6